------------------------------------------------------------ Date: Mon Mar 8 07:46:21 2004, Shift: 00:00-08:00 Supervisor: EBJ, with: Log book#15 pages 79 - 79 Shift Summary: ------------------------------------------------------------ 1. We had a good amount of beam. The machine seemed to be performing fairly well. 2. I took some data runs at the current settings. When they dropped the store around 1am, I took the pedestal runs. 3. There were no detector problems. ------------------------------------------------------------ _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Mon Mar 8 07:46:31 2004
This archive was generated by hypermail 2.1.8 : Mon Mar 08 2004 - 07:46:52 EST